OpenStudy (anonymous):

What is the solution of the following DIFFERENTIAL EQUATION for the initial value problem: y(0) = -2 Eqn: y' + (tanx)y = 2(cosx)^2

OpenStudy (anonymous):

@Mertsj @Zarkon @dpaInc @jim_thompson5910 @lgbasallote @KingGeorge @AccessDenied @SmoothMath @eliassaab @asnaseer @.Sam. @kropot72 @eigenschmeigen @satellite73 @amistre64 @TuringTest @shivam_bhalla @experimentX @FoolForMath @Callisto @phi

OpenStudy (anonymous):

I know integrating factors now. So I get the integrating factor to be: secx.

OpenStudy (anonymous):

After doing the simplification

OpenStudy (anonymous):

But what I am getting when I solve for y is: \[y = 2\cos\]

OpenStudy (anonymous):

Oops

OpenStudy (anonymous):

\[y = 2cosxsinx + C\]

OpenStudy (anonymous):

I know: \[\mu(x) = \frac{1}{cosx}\] mu is the INTEGRATING FACTOR

OpenStudy (anonymous):

I think that \[ y=2 (\sin (x) \cos (x)-\cos (x) \]

OpenStudy (anonymous):

This is after simplification

OpenStudy (anonymous):

But how did you get that -cosx???

OpenStudy (anonymous):

Ok this is what I did:

OpenStudy (anonymous):

I start off with the equation: y' + (tanx)y = 2(cosx)^2 I multiply both sides by the integrating factor. Let's call it a(x), not mu(x) a(x)[y' + tanx*y] = 2(cosx)^2

OpenStudy (anonymous):

Then I expanded:

OpenStudy (anonymous):

a(x)*y' + a(x)*tanx*y = a(x)*(2cosx)^2 <-- I missed that a(x) on the RHS, sorry about that

OpenStudy (anonymous):

Then, we can re-express a(x)tanx: Recall: a(x) = \[e^{\int\limits_{}^{}p(x)dx}\]

OpenStudy (anonymous):

My p(x) is tanx

OpenStudy (anonymous):

The problem is that on the RHS: I have: \[a(x)*2(cosx)^2 \]

OpenStudy (anonymous):

Which will then simplify to: \[secx*(2)(cosx)\]

OpenStudy (anonymous):

cos^2x sorry

OpenStudy (anonymous):

So then you end up getting cosx on the RHS

OpenStudy (anonymous):

Could you show me how you did your RHS

OpenStudy (anonymous):

\[ -\cos(x) \] is the solution of the homogenous equation \[ y'(x)+y(x) \tan (x)=0 \]

OpenStudy (anonymous):

\[2*\sec(x)*(\cos(x)^2) = 2\cos(x)\]

OpenStudy (anonymous):

I get the above?

OpenStudy (anonymous):

\[2*\frac{1}{cosx}*cosx*cosx\] bow two cosx cancel to leave 2cosx or did I miss something

OpenStudy (anonymous):

@eliassaab I found out how I lost that other cos(x)

OpenStudy (anonymous):

Now, I am getting: C = -2

OpenStudy (anonymous):

So, the solution for this initial value problem is: y = 2sinxcosx - 2cosx

OpenStudy (anonymous):

I found out that I have to cross-multiply AFTER I put in that "C" to the other side

OpenStudy (anonymous):

Would this be correct?

OpenStudy (anonymous):

With the -2 in front of the cosx?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes! Finally.

OpenStudy (anonymous):

Finished.